OpenStudy (anonymous):

You should be able to solve this? x - 7 = 14 + x

hero (hero):

You can't solve it unless it has a solution.

OpenStudy (inkyvoyd):

^tautology.

OpenStudy (inkyvoyd):

Argos, what led you to that conclusion?

OpenStudy (inkyvoyd):

ChmE, what led you to that conclusion?

OpenStudy (anonymous):

it's 21x

hero (hero):

lol, no it's not

OpenStudy (inkyvoyd):

@woahitsmaddie , what led you to that conclusion? Guys, please back up final answers with the work needed to obtain them.

OpenStudy (anonymous):

lol,,, yes it is

hero (hero):

I'm not going to provide any work, because the work would be pointless. Clearly there is no solution.

OpenStudy (inkyvoyd):

Clearly there is.

hero (hero):

Okay @inkyvoyd, if you prefer x - 7 = 14 + x x - x = 14 + 7 0 = 21 <-----False (No solution)

OpenStudy (inkyvoyd):

Exactly. THe solution (to the problem) is that there is no solution to the problem. it is inconsistence LOL

hero (hero):

Saying that no solution is a solution is a outright contradiction

OpenStudy (anonymous):

your problem was x-7=14+x... so you would get the x values to one side, by subtracting or adding, then you would add 7 over because it was originally a negative... finally when you add those two together you get 21x, because the x was to the right of the 14 originally x-7=14+x -7=14x 21x

OpenStudy (inkyvoyd):

no, solution and solution are different. I am referring to two different solutions.

OpenStudy (shubhamsrg):

actually, the unreal solution would be x= infinity

OpenStudy (shubhamsrg):

the 2 lines are parallel, the theory says the lines intersect at infinity the above can also be understood if you take the ratio of heights with x->inf ratio comes out to be 1 kudos! ^_^

hero (hero):

No bro, lines extend infinitely and if they are parallel, they never intersect. There's no such thing as lines connecting at infinity. The only way you can say that is if you've been to infinity. No one has, so no one can make such a claim.

OpenStudy (shubhamsrg):

as i said, thats just a theory what i mean is lim x-> inf (x-7)/(14+x) thats ofcorse 1 limits is all theoretical !

OpenStudy (shubhamsrg):

suppose we are checking continuity at x->0 we can not reach the next number to 0 ofcorse, but still, it can be theoretically calculated.

OpenStudy (anonymous):

strange question but, a solution is a value such that the equation is true. the solution is the null set or just what @hero said, no solution

OpenStudy (abb0t):

lol. limits r so random.